Calculate the numbers of electron in 18 ml of water at 4 degree celsius and 1 atm pressure
 

Grade:9

2 Answers

Arun
4 years ago

We know density of water is 1 gm/cc or 1 gm/ml. So 18 ml of water contains 18 gm of water implies 1 mole of water.

By Avogadro no. 1 mole of water contains 6.023 x 10^(23) molecules of water.

But each molecule of water contains 10 electrons.

Thus 6.023 x 10^(23) molecules of water will contain 10 x 6.023 x 10^(23) electrons= 6.023 x 10^(24) electrons

Khimraj
4 years ago
Given, volume of water = 18ml and density of water at 4°C = 1g/ml 
so, mass of water = 18 × 1 = 18g 
Mole of water = mass/molar mass 
= 18g/18g/mol = 1 mol 
∴ number of water molecule = mole of water × Avogadro's number
= 1 × 6.023 × 10²³
= 6.023 × 10²³
We know, there are 10 electrons in a molecule of water .
e.g., 8 electrons of Oxygen atom and 2 electrons of each two hydrogen atoms.
Now, number of electrons = number of water molecule × 10
= 6.023 × 10²³ × 10 
= 6.023 × 10²⁴